OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (anonymous):

@DisplayError

OpenStudy (displayerror):

We can calculate the slope of the line since we are given two points \[\text{Slope} = m = \frac{y_2 - y_1}{x_2 - x_1}\] Because we now have a slope and a point, we can use the point-slope form of a line: \[y - y_1 = m(x - x_1)\] ^ Just plug in the slope that you calculated and a point (either (4,1) or (5,2) will work). From that, rearrange your equation such that it is in the form \[Ax + By = C\] such that A is a positive integer and B and C are integers.

OpenStudy (anonymous):

which is y?

OpenStudy (displayerror):

You only plug in for y_1 and x_1. y and x are the y and x in the standard form equation.

OpenStudy (anonymous):

could u give me the formula with the equations plugged in?

OpenStudy (anonymous):

@DisplayError

OpenStudy (displayerror):

What is your point (x_1, y_1) and your slope m? I'll plug what you give me into the point-slope equation.

OpenStudy (anonymous):

thnx